OpenStudy (anonymous):

Choose the correct sum of the polynomials (4x3 - 2x - 9) + (2x3 + 5x + 3).

OpenStudy (anonymous):

come on help please

OpenStudy (texaschic101):

4x^3 - 2x - 9 + 2x^3 + 5x + 3 --- combine like terms 4x^3 + 2x^3 = 6x^3 -2x + 5x = -3x -9 + 3 = - 6 put it all together and you end up with : 6x^3 - 3x - 6